1.) Anderson Corporation has purchased a group of assets for

$ 20 comma 500$20,500.

The assets and their relative market values are listed below.

Land

$7,300

Equipment

3,000

Building

3,200

Which of the following amounts would be debited to the Land​ account? (Round any intermediate calculations to two decimal​ places, and your final answer to the nearest​ dollar.)

2.)The Golden Company issues $520,000 o f77​%,10−year bonds at 108 on March​ 31, 2019. The bonds pay interest on March 31 and September 30. Assume that the company uses the straight- line method for amortization. The journal entry to record the issuance includes a​ ________.
Answer #1

1) Total market value = 7300+3000+3200 = 13500

Amount allocated to Land account = 7300/13500 = 54.07% = 20500*54.07% = 11084

2) Journal entry

Date account and explanation Debit Credit
Cash (520000*1.08) 561600
Bonds payable 520000
Premium on bonds payable 41600